OpenStudy (anonymous):

Solve the radical equation: the square root of 4+3x + the square root of x+2 =2

OpenStudy (anonymous):

\[\sqrt{4+3x} + \sqrt{x+2} = 2\]

OpenStudy (nikvist):

x=-1